OpenStudy (anonymous):

23 degrees 20' 48" is the same as _____

OpenStudy (stormfire1):

23.346667 degrees.

OpenStudy (anonymous):

\[23+\frac{20}{60}+\frac{48}{3600}=\frac{1751}{75}=23.346667{}^{\circ} \]